OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (3, 4) and has a slope of 2?

OpenStudy (gorv):

y=mx+c is slope intercept form y=2x+c now plug in x and y to get c

OpenStudy (anonymous):

Is c 3 or 4? Or maybe 7 together?

OpenStudy (anonymous):

put x=3 and y=4 in y=2x+c you get 4=3*2+c c=-2 therefore general equation y=2x-2